# Do Not Worry if You Do Not Have Data: Building Pretrained Language Models Using Translationese

### v2

@@ 5.3 Downstream Tasks and Evaluation
We finetune the mini-1k and base-1k models for various classification, regression, and generation tasks. We do some hyperparameter tuning for each task and then repeat them for different data splits. More hyperparameters and evaluation details can be found in Appendix B. For evaluations, we report our primary scores on IndicGLUE Kakwani et al. (2020) and IndicXNLI (iXNLI) Aggarwal et al. (2022) for Hindi and Gujarati and use the validation set of GLUE benchmark Wang et al. (2018) for English. We also experiment with other generation tasks like CNN-Dailymail Nallapati et al. (2016), DailogSum Chen et al. (2021), XL-Sum Hasan et al. (2021), IndicNLG [^9] We only take first 4k examples of IndicNLG test split for each task due the large test split of IndicNLG Kumar et al. (2022), FLoRes-200 Team et al. (2022), IN22-Conv & IN22-Gen Gala et al. (2023) and use standard evaluation metrics suitable for each task like accuracy, f1-score, Rouge-L Lin (2004) and chrF++ Popović (2017).
@@ 6.1 Main Results
Mono syn-EN_hi-filtered + 10% 90.13 35.75 86.41 84.75 84.21 65.34 76.99 76.91 81.95 75.83
@@
Bi BI-EN_syn-HI_syn-filtered + 10% 89.1 35.45 85.34 84.53 84.18 65.7 76.64 77.24 82.1 75.59
In this section, we present results for Hindi, Gujarati, and English language models trained on clean data, as well as synthetic data generated from translations. We demonstrate the impact of filtering and adding additional clean data for extended pretraining of LMs trained solely on synthetic text. Additionally, we observe the effect of using the clean source text along with its translations (synthetic parallel documents) on downstream tasks. We follow the naming convention for different data splits as specified in Section 5. We provide details for pretraining of each model in Appendix B.
Filtered Synthetic Data is Competitive with Web Scraped Data: The results in Table LABEL:tab:hi-gu-main and 1 indicate that syn-HI_en-unfiltered, syn-GU_en-unfiltered, and syn-EN_hi-unfiltered exhibit lower downstream performance compared to their filtered counterparts: syn-HI_en-filtered, syn-GU_en-filtered, and syn-EN_hi-filtered, respectively. It is evident that filtering the synthetic documents using TinyLMs significantly improves the performance of both NLU and NLG tasks. In Table 1, we observe that for tasks like CoLA Warstadt et al. (2019), language models trained solely on synthetic data lag behind when compared to other tasks. This suggests that synthetic corpora may lack certain important elements necessary for language models to perform competitively in linguistic acceptability tasks, as opposed to LMs trained on clean, non-synthetic corpora.
Fine-tuning on Web Scraped Data boosts performance: Even after filtering, we observe that language models trained solely on synthetic text slightly underperform LMs trained on clean text. To address this issue, we conduct extended pretraining of LMs using clean data sourced from IndicMonoDoc. The objective is to determine if this additional training improves performance. We only incorporate an additional 10% of clean data compared to the LM’s previous training data. We see these results across all three languages, and for Hindi and Gujarati, we see that by incorporating even a small amount of clean data, we observe an increase in performance on downstream tasks, bringing the LM at par or closer to the performance of a clean LM. We see an improvement in LMs trained using unfiltered synthetic corpora as well but we believe that filtering leads to the removal of noisy documents and thus better performance.
Model iXNLI bbc-a iitp-mr iitp-pr midas Avg.
HI-clean 68.74 80.25 67.74 77.05 78.33 74.42
syn-HI_en-unfiltered 67.32 77.92 65.63 76.81 77.58 73.05
syn-HI_en-filtered 69.48 78.98 65.16 77.43 77.33 73.68
syn-HI_en-filtered+1070.15 79.56 67.09 78.2 79.03 74.81
Using synthetic for one language doesn’t impact performance in another: For many multilingual language models, data imbalance causes a gap in performance across languages. But what if we can combine synthetic data along with clean data for training multilingual models, would the synthetic part deteriorate the performance of the multilingual model? To experiment with this, we train bilingual base-1k models over different combinations of clean and synthetic corpora for English and Hindi and evaluate their performance on GLUE Wang et al. (2018), and report performance on IndicNLG, and Machine translation in Appendix A. Following Table 1, we see that using Hindi synthetic data does not affect its performance compared to BI-EN-HI-clean model which is solely trained on clean corpora. This implies that it is possible to train multilingual models where some languages are trained only over a clean subset and others on synthetic without deteriorating performance across languages. We further see that using parallel data does not have much impact on multilingual models.
6.2 Further Exploration
sst2 cola mrpc qnli qqp rte mnli-m mnli-mm stsb
Model acc mcc f1 acc f1 acc acc acc pearson Avg.
Translationese Hi->En syn-EN_hi-filtered + 10% 87.84 26.61 83.27 78.5 80.36 61.37 71.29 71.11 71.91 70.25
Translationese Gu->En syn-EN_gu-filtered + 10% 86.58 25.17 81.67 67.1 77.75 57.76 68.78 68.56 27.54 62.32
[…]
Impact of model size: Following Table 3 and 2, we see that even after scaling down we see consistent improvements for filtering and adding additional data, which empirically shows that indeed using synthetic text after filtering is a viable option for pretraining LMs of varying sizes. In Table 2 we see that after filtering and extended pretraining, synthetic text outperforms LMs trained on clean documents from the web in Hindi.
Model XLSum HG XLSum QG Cnn Dialogsum Avg.
EN-clean 23.87 24.05 16.08 20.39 21.10
syn-EN_hi-unfiltered 22.17 22.97 12.56 18.30 19.00
syn-EN_hi-filtered 23.27 23.83 15.88 19.83 20.70
@@ 6.2 Further Exploration
7 Conclusion
In this paper, we performed a first of its kind study of the feasibility of using translationese data for training language models. We proposed a simple pipeline involving the translation of documents at scale followed by filtering using small and efficient language models trained on clean data. We then showed on a variety of downstream natural language understanding and generative tasks that language models trained on unclean synthetic data were only slightly inferior to those trained on original data, however, filtered synthetic data with extended pretraining on clean data mostly eliminates this gap. We also observed a positive impact of synthetic data on TinyLMs fine-tuned on 10% clean data. While we observed that the source language, and potential content, for synthetic data generation matters, it is clear that synthetic data can help bridge the resource scarcity faced by a vast majority of languages for language modeling. As a part of this work, we also created IndicMonoDoc, the largest collection of clean document-level datasets for 22 Indic languages and English, which we release along with our synthetic data, pipelines, and code. In the future, we aim to first generate synthetic data at much larger scales and experiment with large language models to push the boundaries of language modeling for low-resource languages.
Limitations
We consider the following limitations of our work.
  • Work mainly focuses on TinyLMs so not all observations may carry over to large language models, however, synthetic data generated from translations can surely help fill knowledge gaps.
  • We could not experiment with entire test sets of IndicNLG tasks like Question Generation, WikiBio generation, Headline Generation, and Sentence Summarization due to its vast test split but we do not expect the main trends to change given that we already use 4000 examples per language.
  • For GLUE tasks we report our numbers on the validation set and not on the test set for all models since the scale of our experiments was large, automatically submitting results on the test set was not feasible. We follow existing works doing the same. Plus our goal is not to achieve state-of-the-art results but rather to establish the utility of synthetic data by observing trends.
  • We have not manually verified synthetic data so despite cleaning using TinyLMs there are chances that there may be some toxic content or bad documents. However, this is a future work.
  • Our framework places significant emphasis on the translation model’s performance. Nevertheless, we are confident that this approach will significantly contribute to enhancing the performance of mid-resource languages, particularly those for which the translation model demonstrates considerable proficiency.
Ethical Considerations
We aim to release the code and models with an MIT License [^10] https://opensource.org/license/mit/. The dataset will be released under a CC-0 License [^11] https://creativecommons.org/share-your-work/public-domain/cc0/.
Appendix A Additional results
We report additional results in this section. Tables 6, 7 show the chrF++ and BLEU scores across three translation evaluation benchmarks. This shows that using parallel synthetic data does not deteriorate the performance of the language model. Similar results are shown in Table 8 for IndicNLG tasks where performance on Hindi generation tasks are only affected by a small margin and coupled with results in Table 1 showing that scores are not affected by using Hindi synthetic parallel data.
@@ Appendix B Training and Evaluation
B.1 Training
In this section, we list the dataset and hyperparameters used for training our models for the experiments. For the pretraining of the base models, we keep a hard limit for the base-1k model as 2.38B tokens and for the mini-1k model as 1B tokens. But for the TinyLM we relax this token limit until we see overfitting. For our experiments, we use the NVIDIA A100-SXM4-80GB GPUs.

### v3

@@ 4.3 Downstream Tasks and Evaluation
We finetune the mini-1k and base-1k models for classification, regression, and generation tasks. Hyperparameter tuning is performed using the validation set for models trained only with clean data, and this process is repeated for different data splits. More details on hyperparameters and evaluation can be found in Appendix B. Primary scores are reported on IndicGLUE Kakwani et al. (2020a) and IndicXNLI (iXNLI) Aggarwal et al. (2022) for Hindi and Gujarati, and the GLUE benchmark validation set Wang et al. (2018) for English. We also experiment with other generation tasks like CNN-Dailymail Nallapati et al. (2016), DailogSum Chen et al. (2021), XL-Sum Hasan et al. (2021), IndicNLG Kumar et al. (2022), FLoRes-200 Team et al. (2022), IN22-Conv & IN22-Gen Gala et al. (2023) and use standard evaluation metrics suitable for each task like accuracy, f1-score, Rouge-L Lin (2004) and chrF++ Popović (2017). Further details about each of the evaluation datasets can be found in Appendix B.1.
@@ 5.1 Main Results
In this section, we present results for Hindi, Gujarati, and English language models trained on clean data, as well as synthetic data generated from translations. We demonstrate the impact of filtering and adding additional clean data for extended pretraining of LMs trained solely on synthetic text. Additionally, we observe the effect of using the clean source text along with its translations (synthetic parallel documents) on downstream tasks. We follow the naming convention for different data splits as specified in Section 4. We provide details for the pretraining of each model in Appendix B. We provide additional results in Appendix A.
Filtered Synthetic Data is Competitive with Web Scraped Data: The results in Table LABEL:tab:hi-gu-main and 2 indicate that syn-HI_en-unfiltered, syn-GU_en-unfiltered, and syn-EN_hi-unfiltered exhibit lower downstream performance compared to their filtered counterparts: syn-HI_en-filtered, syn-GU_en-filtered, and syn-EN_hi-filtered, respectively. It is evident that filtering the synthetic documents using TinyLMs significantly improves the performance of both NLU and NLG tasks. We also observe that for tasks like CoLA Warstadt et al. (2019), language models trained solely on synthetic data lag behind when compared to other tasks as seen in Table 8 of Appendix A. This suggests that synthetic corpora may lack certain important elements necessary for language models to perform competitively in linguistic acceptability tasks, as opposed to LMs trained on clean, non-synthetic corpora. Results for base-1k for English are presented in Table 8 in Appendix A because we focus our attention on Indic languages.
sst2 cola mrpc qnli qqp rte mnli-m mnli-mm stsb
Model acc mcc f1 acc f1 acc acc acc pearson Avg.
[…]
Fine-tuning on small amounts of Web Scraped Data boosts performance: Even after filtering, we observe that language models trained solely on synthetic text slightly underperform LMs trained on clean text. To address this issue, we conduct extended pretraining of LMs using clean data sourced from IndicMonoDoc. The objective is to determine if this additional training improves performance. We only incorporate an additional 10% of clean data compared to the LM’s previous training data. We see these results across all three languages, and for Hindi and Gujarati, we see that by incorporating even a small amount of clean data, we observe an increase in performance on downstream tasks, bringing the LM at par or closer to the performance of a clean LM. We see an improvement in LMs trained using unfiltered synthetic corpora as well but we believe that filtering leads to the removal of noisy documents and thus better performance. We observe improved performance in language models (LMs) trained with unfiltered synthetic corpora, but filtering out noisy documents enhances performance further. Our ablation study (Table 13 in Appendix A) investigates whether adding 10% more synthetic data contributes to this improvement or if the data type is key. While performance gains could stem from statistical variances, the consistency across nearly all downstream tasks suggests otherwise.
[…]
Model iXNLI bbc-a iitp-mr iitp-pr midas Avg.
HI-clean 68.74 80.25 67.74 77.05 78.33 74.42
syn-HI_en-unfiltered 67.32 77.92 65.63 76.81 77.58 73.05
syn-HI_en-filtered 69.48 78.98 65.16 77.43 77.33 73.68
syn-HI_en-filtered+1070.15 79.56 67.09 78.2 79.03 74.81
5.2 Further Exploration
Analysis of Synthetic Data: Figure LABEL:fig:_variance_plot shows the perplexity mean and variance scores for TinyLM across token positions in the documents. This shows that on unseen documents, TinyLM shows higher variance on English documents generated by translating Gujarati documents from IndicMonoDoc as compared to English clean and English synthetic generated from Hindi. This also gives reason for the deterioration in results in Table 2 due to Gujarati documents. Figure 7 shows the distribution of lengths of filtered documents by TinyLMs showing that they do not add any bias for shorter documents during filtering.
Impact of model size: Following Table 2 and 3, we see that even after scaling down we see consistent improvements for filtering and adding additional data, which empirically shows that indeed using synthetic text after filtering is a viable option for pretraining LMs of varying sizes. In Table 3 we see that after filtering and extended pretraining, synthetic text outperforms LMs trained on clean documents from the web in Hindi. This is also supported by our experiments on finetuning Llama-3-8B in Section 5.3.
Model XLSum HG XLSum QG Cnn Dialogsum Avg.
EN-clean 23.87 24.05 16.08 20.39 21.10
syn-EN_hi-unfiltered 22.17 22.97 12.56 18.30 19.00
syn-EN_hi-filtered 23.27 23.83 15.88 19.83 20.70
5.3 Scaling to Llama-3-8B
To show the effect of using translationese on larger models, we select Llama-3-8B [^12] https://github.com/meta-llama/llama3 and Gemma-2B (Team et al., 2024) and perform continual pretraining over clean and synthetic data to improve ability over the low resource target language. We take Marathi as a replacement for Hindi for scaling experiments since data for Hindi is abundantly available and existing models already have a good language understanding of Hindi making it harder to compare the effects of utilizing clean vs synthetic data. For a fair comparison, we limit each data split to 344M tokens for Gujarati and 465M tokens for Marathi and follow a similar procedure as described in Section 3.4 to generate and filter data for Marathi. We perform extended training for a single epoch using LoRA (Hu et al., 2021) finetuning on $W_{q}$, $W_{v}$ projection matrices using $\alpha$=$16$ and $r$=$8$. We keep the learning rate at $3e^{-5}$ with a weight decay of $0.01$ and an effective batch size of 58k.
Perplexity: We report average sentence level perplexity on sentences from IN22-Conv and IN22-Gen Gala et al. (2023) in Table 5. We see that filtered synthetic data for Gujarati outperforms clean data, but for Marathi, it does not. This means that filtering improves performance at scale but relies on the quality of translationese in the target language. We report perplexities on individual test sets in Appendix A.
Data Marathi Gujarati
Gemma
2B
Llama-3
8B
Gemma
2B
Llama-3
8B
Base model 178.898 66.740 71.136 2.839
clean 37.599 11.196 10.350 2.312
synthetic-unfiltered 92.813 15.697 10.941 2.816
synthetic-filtered 104.148 14.622 10.150 2.236
Table 5: Average perplexity ($\downarrow$) of models trained on Translationese vs. Clean data on IN22-Conv and IN22-Gen test sets shows improvement with large-scale models. Bold represents best among synthetic data splits.
[…]
Data Flores-200 IN22-Gen
Marathi Gujarati Marathi Gujarati
Base model 27.83 34.35 30 34.92
clean 34.02 35.61 33.94 35.42
synthetic-unfiltered 30.63 34.19 29.67 32.1
synthetic-filtered 31.81 35.54 31.3 35
Table 6: chrF++ scores on 5-shot Machine Translation FloRes and IN22-Gen test sets on Llama-3-8B. Bold represents best among synthetic data experiments.
Data Marathi Gujarati
Gemma
2B
Llama-3
8B
Gemma
2B
Llama-3
8B
Base model 90.89 ±0.005 95 ±0.009 83.84 ±0.002 92.69 ±0.018
clean 90 ±0.014 97.17 ±0.013 87.79 ±0.012 93.33 ±0.0132
synthetic
unfiltered
89.66 ±0.012 95.38 ±0.016 83.88 ±0.009 92.81 ±0.01
synthetic
filtered
86.67 ±0.016 96.15 ±0.011 84.10 ±0.013 92.94 ±0.007
Table 7: Average accuracy with standard deviation (superscript) over 5 runs on 10-shot IndicSentiment classification task. Bold represents the best among synthetic data experiments.
6 Conclusion
In this paper, we performed a first of its kind study showing the promise of using translationese data for training language models for low-resource languages. Our simple pipeline involves the translation of high-resource source language documents at scale, followed by perplexity based filtering using small and efficient language models trained on clean target low-resource language data. We then showed on a variety of downstream natural language understanding and generative tasks that both small and large language models pre-trained on clean synthetic data are comparable to those trained on clean data. While we observed that the source language for synthetic data generation matters, it is clear that synthetic data can help bridge the resource scarcity faced by a vast majority of languages for language modeling. Future work will focus on better and faster synthetic data generation and filtering mechanisms.
Acknowledgements
We extend our thanks to the Department of Computer Science and Engineering at the Indian Institute of Technology (IIT) Bombay for providing access to GPU servers, and to the Centre for Development of Advanced Computing (C-DAC) for granting us access to the Param Siddhi Supercomputer. These computational resources were essential for the successful completion of this work.
Limitations
We consider the following limitations of our work.
  • We show that synthetic data also helps for larger models like Llama-3-8B but for even larger models above 100B parameters, effects of translationese may be different. However, synthetic data generated from translations can surely help fill knowledge gaps.
  • Due to the extensive size of the test sets for IndicNLG tasks (Question Generation, WikiBio generation, Headline Generation, and Sentence Summarization), we couldn’t experiment with them in their entirety. However, since we already use 4000 examples per language, we anticipate that the overall trends remain unchanged.
  • We report GLUE validation set results for all models due to the large scale of our experiments, following existing practices. Our goal is to demonstrate synthetic data utility, not to achieve state-of-the-art results.
  • Our framework heavily relies on the translation model’s performance. Despite this dependency, we are confident that our approach will significantly enhance the performance of mid-resource languages, especially where the translation model is already proficient.
Ethical Considerations
We release the code and models with an MIT License [^13] https://opensource.org/license/mit/. The dataset is released under a CC-0 License [^14] https://creativecommons.org/share-your-work/public-domain/cc0/.
Appendix A Additional results
We report additional results in this section. Tables 14, 15 show the chrF++ and BLEU scores across three translation evaluation benchmarks. This shows that using parallel synthetic data does not deteriorate the performance of the language model. Similar results are shown in Table 16 for IndicNLG tasks where performance on Hindi generation tasks are only affected by a small margin and coupled with results in Table 8 showing that scores are not affected by using Hindi synthetic parallel data.
@@ Appendix A Additional results
syn-EN_hi-filtered + 10% 90.13 35.75 86.41 84.75 84.21 65.34 76.99 76.91 81.95 75.83
Using synthetic for one language doesn’t impact performance in another: For many multilingual language models, data imbalance causes a gap in performance across languages. But what if we can combine synthetic data along with clean data for training multilingual models? would the synthetic part deteriorate the performance of the multilingual model? To experiment with this, we train bilingual base-1k models over different combinations of clean and synthetic corpora for English and Hindi and evaluate their performance on GLUE Wang et al. (2018), and report performance on IndicNLG, and Machine translation in Appendix A. Following Table 9, we see that using Hindi synthetic data does not affect its performance compared to BI-EN-HI-clean model which is solely trained on clean corpora. This implies that it is possible to train multilingual models where some languages are trained only over a clean subset and others on synthetic without deteriorating performance across languages. We further see that using parallel data does not have much impact on multilingual models.
sst2 cola mrpc qnli qqp rte mnli-m mnli-mm stsb
Model acc mcc f1 acc f1 acc acc acc pearson Avg.
@@
BI-EN_syn-HI_syn-filtered + 10% 89.1 35.45 85.34 84.53 84.18 65.7 76.64 77.24 82.1 75.59
Table 9: Results on English for Bilingual models: Dev set of GLUE tasks for different synthetic splits on the base-1k model. Training bilingual models using synthetic data in one language (Hindi) does not affect the performance in the other language (English). Bold values represent the best amongst synthetic splits.
@@
Perplexity-IN22 Conv
Data Marathi Gujarati
Gemma-2B Llama-3-8B Gemma-2B Llama-3-8B
Base model 332.9036 121.6586 131.8027 3.3922
clean 58.3804 15.1342 12.3128 2.5208
synthetic-unfiltered 168.0933 23.018 12.7995 3.0671
synthetic-filtered 191.1506 21.355 11.6916 2.3849
Table 11: Perplexity on IN-22 Conv. Bold values represent best among synthetic splits.
Perplexity-IN22 Gen
Data Marathi Gujarati
Gemma-2B Llama-3-8B Gemma-2B Llama-3-8B
Base model 24.893 11.821 10.4693 2.2853
clean 16.8172 7.2572 8.3868 2.1032
synthetic-unfiltered 17.5333 8.3766 9.0821 2.564
synthetic-filtered 17.1451 7.8885 8.6089 2.0868
Table 12: Perplexity on IN-22 Gen. Bold values represent best among synthetic splits.
NLU NLG
Model iXNLI bbc-a iitp-mr iitp-pr midas Avg.
Headline
Gen.
Sentence
Summ.
Question
Gen.
Wikibio Avg.
HI-clean 73.61 81.75 72.58 79.73 80.34 77.60 27.54 23.64 24.84 52.16 32.04
syn-HI_en-unfiltered 72.87 77.92 64.36 76.22 79.91 74.26 27.29 22.93 24.22 50.14 31.14
syn-HI_en-unfiltered+10clean 74.63 78.36 67.75 77.46 80.17 75.67 26.98 23.20 24.76 51.34 31.57
syn-HI_en-filtered 74.75 81.06 69.03 78.58 79.73 76.63 27.15 23.10 24.41 49.88 31.13
syn-HI_en-filtered+10clean 74.49 80.94 71.61 79.92 80.64 77.52 27.87 24.23 24.87 51.18 32.04
syn-HI_en-filtered+10synthetic 74.95 80.83 71.93 79.92 78.73 77.27 27.64 25.13 23.60 43.48 29.96
Table 13: Ablation Results for Hindi using additional 10% clean vs synthetic data: NLU/NLG tasks on base-1k (Hindi) different clean and synthetic splits. Test accuracy for NLU tasks; Rouge-L F1 scores for NLG tasks. Bold values represent the best amongst synthetic splits.
@@
IndicXNLI 10-Shot
Data Marathi Gujarati
Gemma-2B Llama-3-8B Gemma-2B Llama-3-8B
Base model 36.56 ± 0.47 50.66 ± 1.46 34.74 ± 0.83 47.63 ± 1.96
clean 35.92 ± 0.79 48.47 ± 1.44 33.40 ± 0.29 48.01 ± 0.99
synthetic-unfiltered 35.66 ± 0.93 49.50 ± 1.76 33.26 ± 0.39 48.22 ± 1.88
synthetic-filtered 34.4 ± 0.23 51.98 ± 1.32 33.78 ± 0.22 47.41 ± 2.17
Table 17: Classification Results on IndicXNLI using 10-shot prompting.
@@ Appendix B Training and Evaluation
In this section, we provide an overview of the training and evaluation setup employed in our experiments. This includes details about the datasets used, training hyperparameters, evaluation metrics, and other relevant configurations.
B.1 Evaluation Datasets
For evaluation, we utilize a diverse set of datasets covering four languages: English, Hindi, Gujarati, and Marathi. For Hindi and Gujarati, we rely on the IndicGLUE benchmark [^15] https://huggingface.co/datasets/ai4bharat/indic_glue Kakwani et al. (2020b), which provides a range of tasks for natural language understanding (NLU), including natural language inference (IndicXNLI/iXNLI), article genre classification (bbc-a, iNLTK), discourse mode classification (MIDAS), and sentiment analysis (iitp-mr, iitp-pr). For natural language generation (NLG), we employ the IndicNLG benchmark [^16] https://huggingface.co/collections/ai4bharat/indicnlg-66c5a1397bab135be074cfe1 Kumar et al. (2022), which includes tasks like headline generation, sentence summarization, question generation, and Wikipedia biography generation.
The IndicGLUE dataset is semi-automatically curated using website metadata and Wikipedia articles, while IndicNLG is derived from Wikipedia articles and news websites for summarization, along with parallel corpora and pivot-based translation for paraphrasing tasks. Additionally, we incorporate the test sets from IN22 Gala et al. (2023) and Flores-200 Team et al. (2022) to evaluate performance on machine translation tasks.
We use the well-known GLUE benchmark, which includes nine NLU tasks in English such as natural language inference (NLI), semantic similarity, text classification, and linguistic acceptability. For English summarization tasks, we rely on XLSum Hasan et al. (2021), DialogSum Chen et al. (2021), and CNN/DailyMail See et al. (2017).
B.2 Training
For the pretraining of the base models, we keep a hard limit for the base-1k model as 2.38B tokens and for the mini-1k model as 1B tokens. But for the TinyLM we relax this token limit until we see overfitting. For our experiments, we use the NVIDIA A100-SXM4-80GB GPUs.